Suppose that a competitive firm's marginal cost of producing output q is given byMC(q) = 3+2q. Assume that the market price of the firm's product is $9. (a) What level of output will the firm produce?

(b) ]Suppose that the average variable cost of this rm is given by AV C(q) = 3 + q. Suppose that the firm's fixed costs are known to be $3. Will the firm be earning a positive, negative, or zero prot in the short run? Show your result on a graph.
Answer #1

A) at eqm, P = MC

9 = 3 + 2q

q* = 6/2 = 3

b) AVC = 3+q

FC = 3

So AC = 3+q+3/q

So in short run,

At q= 3, AC = 3+3+1 = 7

So as P > AC

So firm will earn positive profit in short run

graph

green colour rectangle is profit